# On Approximate Representation of Fractional Brownian Motion

### v2

@@
00footnotetext: Email: rkoffice@mail.ru
@@ 5. Approximate Representation of Fractional Brownian Motion
The mean square approximation errors $\varepsilon$ and $\varepsilon^{*}$ are given in Tables 1 and 2 for $T=1$, $L=4,8,\dots,128$, and $H=0.1,0.2,\dots,0.9$.
@@
$H$ $L=4$ $L=8$ $L=16$ $L=32$ $L=64$ $L=128$
0.1 0.385941 0.324399 0.273198 0.230844 0.195706 0.166421
0.2 0.186654 0.136295 0.100453 0.074599 0.055705 0.041760
0.3 0.103540 0.065560 0.042260 0.027516 0.018017 0.011834
0.4 0.060718 0.033052 0.018491 0.010483 0.005981 0.003424
0.5 0.035714 0.016667 0.008065 0.003968 0.001969 0.000980
0.6 0.020484 0.008212 0.003436 0.001466 0.000632 0.000274
0.7 0.013274 0.004949 0.001927 0.000764 0.000305 0.000123
0.8 0.021554 0.011519 0.006396 0.003602 0.002043 0.001166
0.9 0.081270 0.061755 0.046945 0.035625 0.027012 0.020475
On a qualitative level, these results are close to those published in [3]. Here, we mean the convergence rate depending on $H$ that is defined by the analysis of the above tables, although another accuracy measure of the approximation was used in [3]. However, it is interesting to note that even for small $L$, the data from Tables 1 and 2 are very close. This implies that the representation of the matrix $K^{H}$ as a product of four matrices has almost no effect on the accuracy for the approximate representation but allows one to form computationally stable algorithms (the numerical results presented in this paper were obtained using symbolic calculations in Mathcad computer algebra system). For Hurst indices close to the boundaries of the interval $(0,1)$, the high relative error and low convergence rate are largely due to properties of the function $k_{H}(\cdot)$. To calculate the relative error, it is sufficient to compare the data from Tables 1 and 2 with the squared norm of the function $k_{H}(\cdot)$, and this squared norm is equal to $1/(2H+1)$ for $T=1$.
@@
The mean square approximation error $\varepsilon^{*}$ can be calculated exactly by the expression proved in Theorem 2. The results obtained by Equation (27) for $H=0.6,0.7,0.8,0.9$, as well as the results obtained by Equation (28) for $H=0.1,0.2,0.3,0.4$, are given in Table 3 for $T=1$ and $L=4,8,\dots,128$. For $H=1/2$, Equations (27) and (28) provide the same result. It is easy to see that the data from Tables 2 and 3 differ insignificantly.

### v3

@@
00footnotetext: Email: rkoffice@mail.ru00footnotetext: Citation: Rybakov, K.A. On Approximate Representation of Fractional Brownian Motion. Methodol. Comput. Appl. Probab. 2025, 27(4), 88. https://doi.org/10.1007/s11009-025-10219-w
@@ 5. Approximate Representation of Fractional Brownian Motion
The mean square approximation errors $\varepsilon$ and $\varepsilon^{*}$ are given in Tables 1 and 2 for $T=1$, $L=4,8,\dots,1024$, and $H=0.1,0.2,\dots,0.9$.
@@
$H$ $L=4$ $L=8$ $L=16$ $L=32$ $L=64$ $L=128$ $L=256$ $L=512$ $L=1024$
0.1 0.385941 0.324399 0.273198 0.230844 0.195706 0.166421 0.141905 0.121296 0.103905
0.2 0.186654 0.136295 0.100453 0.074599 0.055705 0.041761 0.031396 0.023653 0.017846
0.3 0.103540 0.065560 0.042260 0.027516 0.018017 0.011835 0.007788 0.005130 0.003381
0.4 0.060718 0.033052 0.018491 0.010483 0.005981 0.003424 0.001963 0.001127 0.000647
0.5 0.035714 0.016667 0.008065 0.003968 0.001969 0.000980 0.000489 0.000244 0.000122
0.6 0.020484 0.008212 0.003436 0.001466 0.000632 0.000273 0.000119 0.000052 0.000022
0.7 0.013274 0.004949 0.001927 0.000764 0.000305 0.000123 0.000050 0.000020 0.000008
0.8 0.021554 0.011519 0.006396 0.003602 0.002043 0.001164 0.000666 0.000381 0.000219
0.9 0.081270 0.061755 0.046945 0.035625 0.027012 0.020475 0.015518 0.011760 0.008913
On a qualitative level, these results are close to those published in [3]. Here, we mean the convergence rate depending on $H$ that is defined by the analysis of the above tables, although another accuracy measure of the approximation was used in [3]. However, it is interesting to note that even for small $L$, the data from Tables 1 and 2 are very close. This implies that the representation of the matrix $K^{H}$ as a product of four matrices has almost no effect on the accuracy for the approximate representation but allows one to form computationally stable algorithms (the numerical results presented in this paper were obtained using C++ program***The source code is available at https://github.com/rkoffice/fBm with Boost.Multiprecision library). For Hurst indices close to the boundaries of the interval $(0,1)$, the high relative error and low convergence rate are largely due to properties of the function $k_{H}(\cdot)$. To calculate the relative error, it is sufficient to compare the data from Tables 1 and 2 with the squared norm of the function $k_{H}(\cdot)$, and this squared norm is equal to $1/(2H+1)$ for $T=1$.
@@
The mean square approximation error $\varepsilon^{*}$ can be calculated exactly by the expression proved in Theorem 2. The results obtained by Equation (27) for $H=0.6,0.7,0.8,0.9$, as well as the results obtained by Equation (28) for $H=0.1,0.2,0.3,0.4$, are given in Table 3 for $T=1$ and $L=4,8,\dots,1024$. For $H=1/2$, Equations (27) and (28) provide the same result. It is easy to see that the data from Tables 2 and 3 differ insignificantly.